A 38.23 g solid sample of mercury is heated from -39 oC its melting point, completely melted then heated to 109 oC, How much energy was absorbed by the sample? (Answer in kJ)

S=0.14 J/goC Delta Hf = 0.011 kJ/g


1
Expert's answer
2021-03-23T04:21:55-0400

Solution:

This problem can be summarized thusly:

1) A sample of solid mercury is melted at -39°C. The heat absorbed is calculated by multiplying the grams of solid mercury by the heat of fusion (ΔHf) in kJ/g.

q1 = ΔHf × m = 0.011 kJ/g × 38.23 g = 0.4205 kJ


2) A sample of liquid mercury is heated from -39°C to 109°C. The heat absorbed is calculated by using the specific heat of liquid mercury (S = 0.14 J/g°C) and the equation: q = m × S × ΔT.

q2 = m × S × (Tf - Ti)

q2 = 38.23 g × 0.14 J/g°C × (109 - (-39))°C = 792.1256 J = 0.7921 kJ


q = q1 + q2 = 0.4205 kJ + 0.7921 kJ = 1.2126 kJ = 1.213 kJ

q = 1.213 kJ


Answer: 1.213 kJ of energy was absorbed by the sample.